Session: Feb 2005 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (K. Borkowski at Tue Mar 8 08:41:18 2005) Minor failures :
BBC3 unlocked (same problem as in GD018). No other known problems.

Medicina (Andrea Orlati at Tue Mar 8 08:59:55 2005) Minor failures : LATE
schedule starts 10 minutes late.

Onsala (Michael Lindqvist at Tue Mar 8 09:13:30 2005) Success :
No known problems.

Jodrell_Bank (Alastair Gunn at Thu Mar 10 22:24:14 2005) Success :
The Lovell Telescope was used. The schedule was altered to limit the number of phase calibrator source changes; some scans will not fringe. No known problems.

Cambridge (Alastair Gunn at Thu Mar 10 22:24:54 2005) Minor failures :
The microwave link failed prior to this experiment due to heavy snow. Data may be good after about 15:48 UT, but link may have been intermittent. No other known problems.

Effelsberg (A. Kraus at Fri Mar 11 12:53:24 2005) Minor failures :
Computer problems (frontend control) in the beginning: 11 mins late. Stopped due to snow: UT1931-2016, 2355-0013, 0428-0440.

Westerbork (Tony Foley at Fri Mar 11 13:47:14 2005) Minor failures :
Wb used in phased array mode. The snow has gone! Tsys getting back to normal

Noto (P. Cassaro at Mon Mar 21 10:10:42 2005) Success :